A 6 L bottle contains 16·1023 N2 molecules. Find the

a) amount of substance and

b) the pressure of the nitrogen gas if its temperature is 0 °C and it is considered ideal.


Could you explain b) in detail please?


1
Expert's answer
2021-06-16T15:38:22-0400

a)

"\nu=\frac{N}{N_A}=\frac{16\cdot 10^{23}}{6.02\cdot 10^{23}}=2.66~mol,"

b)

"p=\frac{\nu RT}{V}=\frac{2.66\cdot 8.31\cdot 273}{0.006}=9.9~atm."